1. 2x^2-x=-6
2. Factor c^2+14c+49=
3. (Use radical exponents to simplify) 4 radical x^12
4. Factor. 144c^2 + 648cf +729f^2=
5. Factor. 6x^7-18x^6+42x^5

Answer by checkley77(12844)

1. 2x^2-x=-6
2x^2-x+6=0
(2x+3)(x-2)=0
2x+3=0
2x=3
x=3/2 ans.
x-2=0
x=2 ans.
-----------------------------------
2. Factor c^2+14c+49
(c+7)^2 ans.
-------------------------------------
3. (Use radical exponents to simplify) 4 radical x^12
4*x^6=4x^6 ans.
---------------------------------------
4. Factor. 144c^2 + 648cf +729f^2=
(12c+27f)^2 ans.
------------------------------------
5. Factor. 6x^7-18x^6+42x^5
6x^5(x^2-3x+7)